40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Apr 6 11:51:15 2010, Koji, Laser, Doubling, MZ lock response.pngoltf.png
    Reply  Wed Apr 7 21:29:48 2010, Dmass, Laser, Doubling, MZ lock 
Message ID: 715     Entry time: Tue Apr 6 11:51:15 2010     Reply to this: 720
Author: Koji 
Type: Laser 
Category: Doubling 
Subject: MZ lock 

Dmass and Koji

Last night we worked on the MZ lock for the doubling experiment.
Dmass made a custom circuit for sum/diff and servoing for the MZ locking.
However it did not lock the MZ when I came into the room.

After some modification of the servo filter, we eventually succeeded to lock it.

Moral of the story:
All of the servo systems should have this kind of response analysis
so that we can easily improve the servo design and can analyze noises from various noise injection points.


1) Measured the PZT response.

The PZT driver has the dewhitening stages with two pairs fo 1Hz single pole / 10Hz single zero. We put 7.3Vpp@0.5Hz to the PZT driver input and saw a complete single fringe wrapping. This means that the PZT response is 0.43 rad/V@0.5Hz. ==> DC gain is 0.54

2) Measured the MZ response.

The amplitude of the MZ fringe was measured at the error point (after the diff amp for the 2 IR-PDs). It was 18.5Vpp. This means the slope at the operating point is 9.25 V/rad.

3) Measured the sevo filter response.

The swept sine signal was injected from the IR PD input, while the TF was measured from the error point to the PZT driving output. It shows +24dB@1Hz and Dmass knew that there is a single pole @0.6Hz. ==> DC gain is 31.


Responses of the components are shown in the first plot. The combined openloop tf is shown in the second plot (BLUE)

It turned out that the gain is too small by a factor of 20~30 despite we only have the room to increase the gain by 5.5.

Thus we decided to move the cut off of the filter from 0.6Hz to 16Hz. Then the servo was expected to be unconditionally stable and the UGF is pushed up ~100Hz, including the increase of the gain by 5.5. The modified openloop tf is shown in the second plot (RED)

Once the servo was modified, it immediately locked. Yeeeeah!

Attachment 1: response.png  45 kB  | Hide | Hide all
response.png
Attachment 2: oltf.png  41 kB  | Hide | Hide all
oltf.png
ELOG V3.1.3-